OpenStudy (anonymous):

Can someone pls help me with these. I need to get the method of solving such questions. Find the complete solution in radians of each equation. a) 2sin^2A + cosA - 1 = 0 b) sin^2A - 1 = cos^2A c) 2sinA + 1 = cscA

OpenStudy (amistre64):

play around with identities to change how the problems look

OpenStudy (anonymous):

In a) should I substitute 1- cos^2A for sin^2A ??

OpenStudy (amistre64):

id give it a shot and see what happens, yes

OpenStudy (anonymous):

The book does not have such sums solved as example but they are given in the exercise....

OpenStudy (amistre64):

these look like they can turn into quadratic equations

OpenStudy (anonymous):

Yes somebody did suggest such a course of action but I have no examples to see , understand and follow..

OpenStudy (amistre64):

2sin^2 + cos - 1 = 0 2(1-cos^2) + cos - 1 = 0 2- 2cos^2 + cos - 1 = 0 - 2cos^2 + cos + 1 = 0 2cos^2 - cos - 1 = 0 cosA = (-b +- sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

can i find some solved examples on some website???

OpenStudy (amistre64):

im sure you can, i just dunno where. if you wanna try some practice problems, try interactmath.com its free and has solutions that you can thumb thru as well

OpenStudy (amistre64):

hit enter, pick a book that matches your material, and play away :)

OpenStudy (anonymous):

Thanks, I'll try the site..

OpenStudy (amistre64):

good luck :)